OpenStudy (anonymous):

Solve: -4(x + 1) - 3 = -3(x - 4)

OpenStudy (anonymous):

distribute first. times the number on the outside of the parenthisis by everything inside them -4(x + 1) - 3 = -3(x - 4) ----> -4x -4 -3 = -3x +12 combine like terms -4x-7 = -3x +12 get one variable by itself, its easier to add in this one, so add the 7. you could subtract the 12 from both sides too if you like -4x = -3x +19 now lets add the -3x to both sides! -x = 19 the x cant be negitive in the answer, so divide by -1 youre answer is x=-19